# The Riemann Hypotheses is false

### v25

@@
say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>4$ be a fixed real number such that $k(1-\Theta-\varepsilon)>2$. For $0<a<b\leq 2a$, note that
@@
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon$. By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (8)
uniformly for $t\geq 2$. Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
@@
$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
Notice that the above integral has a simple pole at $w=1-s$ with residue $\frac{\zeta^{\prime}}{\zeta}(s)-\frac{(T/2\log^{k}T)^{1-s}}{1-s}.$ Hence, shifting the line of integration in (10) to $\Re(s)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
@@
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+iT}\Lambda(m)}{m^{\Theta+\varepsilon+iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (13)
@@
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iT}}{m^{\Theta+\varepsilon+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (14)
uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$. The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation, it also follows from (6) that
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s\ll T^{1/2}\log^{k+1}T.$ (15)
Recall that $H=3T+O(1)$ and $kc>2$. Thus, summing (12) and (13) for $\frac{T}{\log^{k}T}\leq n\leq T$ yields
@@
since $\sum_{m<x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have

### v26

@@
say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>6$ be a fixed real number such that $k(1-\Theta-\varepsilon)\geq 3$. For $0<a<b\leq 2a$, note that
@@
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon/2$. By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t.$ (8)
Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
@@
$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-4iT}^{c+4iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
Notice that the above integrand has simple poles at $w=0,w=\rho$ and $w=1-s$. Hence, shifting the line of integration in (10) to $\Re(w)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+\sum_{2T\leq\gamma_{T}\leq H}\frac{(T/2\log^{k}T)^{\rho-s}}{\rho-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
@@
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\frac{1}{2\pi i}\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{\rho_{T}}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (13)
$+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (14)
By the residue theorem again, we have
$\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ $=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}+$ (15)
$\int_{-\varepsilon+2iT}^{-\varepsilon+iH}\Bigg{)}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s$ (16)
$=\sum_{2T\leq\gamma_{T}\leq H}(T/2\log^{k}T)^{-\rho_{T}}n^{\rho_{T}-1}-\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+iH}^{\Theta+\varepsilon+2iT}+\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\Bigg{)}$ (17)
$\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s+O(n^{-1-\varepsilon}T^{-\varepsilon}).$ (18)
Inserting (18), (17), (16) and (15) into the summation in (13), gives
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+2iT}\Lambda(m)}{m^{\Theta+\varepsilon+2iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (19)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+\sum_{2T\leq\gamma_{T}\leq H}(n^{\rho_{T}-1}+O(n^{-1-\varepsilon}T^{-\varepsilon})).$ (20)
@@
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{2iT}}{m^{\Theta+\varepsilon+2iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (21)
uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$. The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation again, it follows from (7) that
$\sum_{\frac{T}{2\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+2iT}^{-\varepsilon+2iT}\frac{(T/2\log^{k}T)^{-s}n^{s-1}}{\rho_{T}-s}\mathrm{d}s\ll T^{1/2}\log^{k+2}T.$ (22)
By Theorem 4.11 of [6], we have
uniformly for $2T\leq\gamma_{T}\leq H$. Hence
Recall that $H=3T+O(1)$ and $kc>3$. Summing (19) and (20) for $\frac{T}{\log^{k}T}\leq n\leq T$ and invoking the above estimates therefore yields
@@
since $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have